### v1

# A critical assessment of the role of topology on protein thermal stability

Abstract
For several decades, experimental and computational studies have been used to investigate the potential functional role of knots in protein structures. A property that has attracted considerable attention is thermal stability, i.e., the extent to which a protein retains its native conformation and biological activity at high temperatures, without undergoing denaturation or aggregation. Thermal stability is quantified by the melting temperature , an equilibrium property that corresponds to the peak of heat capacity in differential scanning calorimetry experiments. Since protein energy does not depend on the topological state, equilibrium properties such as should not depend on the topology, provided that the exploration of the equilibrium space is ergodic. However, experimental and computational results reported in the literature provide conflicting views on this problem, with some studies reporting an enhancement of thermal stability for certain knotted proteins. Here, we use extensive Monte Carlo simulations of a simple C-alpha protein model of protein YibK to show in a comprehensive manner that does not depend on the topological state of the protein.
1 Introduction
A knotted protein is a protein whose polypeptide chain is self-tied, forming a physical (i.e., open) knot in its native structure. Although the first knotted protein was reported in 1977 richardson1977, it was not until 2000 that Taylor’s computational methodology (which enables a systematic determination of the topological state of open polymer chains) taylor2000 brought these complex proteins into the spotlight. Knotted proteins make up 1% of all protein entries (225 K) in the Protein Data Bank (PDB) Sulkowska2018, and are statistically rare compared to compact random loops of the same size Grosberg2006. All three domains of life contain knotted proteins Virnau2006, including Archaea, a group of single-celled organisms believed to have inhabited the planet for around 3.8 billion years. A recent survey that analyzed 700K knotted proteins predicted by the computer artificial intelligence system AlphaFold (AF) AlphaFold, and 42 proteomes (16 Bacteria, 9 Fungi, 4 Plant, and 13 Animal), reported that the percentage of knotted proteins in any given proteome is around 0.4% Sulkowska2024. The minimum number of crossings in all possible planar projections of a given chain is an incomplete topological invariant used to classify both topological knots (which are closed curves in space) and protein knots. The (or trefoil) knots, a class that features three crossings, is by far the most prevalent protein knot, followed by the and knots, with four and five crossings, respectively Sulkowska2018, Sulkowska2024. The most complex protein knot has seven crossings (it is a knot) and was predicted by AF AlphaFold. It is an 89-residue-long bacterial protein (termed Q9PR55) whose atomic structure was recently experimentally resolved, validating AF’s predictions Danny7_1.
In the context of structural biology, the design principle ‘form follows function’ means that the structure of a protein is adapted to successfully perform its specific task. It highlights the fact that protein structure is not arbitrary, being adjusted by the specificities of its role, which ultimately contributes to the organism’s survival. Additionally, the fact that knots in proteins have been conserved throughout evolution Joanna2012 (and, in many cases, have likely evolved from knotted precursor structures in organisms dating back billions of years) suggests that they may confer some advantage to their carriers. Thus, it is not surprising that a large body of theoretical and experimental work has been devoted to determining the potential functional roles of knots in proteins, alongside establishing how they fold into their intricate native structures. While significant advancements have been made regarding folding and knotting mechanisms (reviewed in Tubiana2024), a consensus has not been reached concerning the function of protein knots (reviewed in Jackson2020, Tubiana2024, Danny2023). Knots can potentially influence proteins at both the physical and structural levels, as well as at the level of their biological function. A recent study found that the knot in protein UCHL-1 is critical for the preservation of secondary structure Sara2024. Previous studies have reported that a knotted structure helps shape and form the binding site of enzymes Nureki2002, Nureki2004, Jacobs2002. Additionally, one study has shown that the knotted region is directly involved in enzymatic activity. Indeed, for the protein TrmD (an example of an important class of enzymes, the SPOUT and SAM synthases), it was shown that the knot regulates catalysis, being critical to protein function Ya-Ming2016. Examples of physical properties are thermal stability, which measures a protein’s resistance to melting and is quantified by the melting temperature; kinetic stability, which measures the protein’s resistance to unfolding at a certain temperature and is quantified by the unfolding rate; and mechanical stability, a measure of the protein’s ability to withstand physical forces and deformations without unfolding or losing its native structure. Simulation studies using models with different levels of resolution suggested that knots in proteins enhance kinetic stability Soler2014, Joanna2008, while an enhancement of mechanical stability has been reported for several model systems based on in vitro experiments Danny2020, Danny7_1(reviewed in Danny2023) and computational studies Joanna2008, Xu2018. Some years ago, Monte Carlo simulations of a simple lattice model were used to find that a knotted topology does not affect thermal stability Soler2013. However, this is in contradiction with a recent experimental study that reported a significant enhancement in thermal stability for a protein featuring a deep trefoil knot Danny2019a. The purpose of the present work is to revisit this question. Our results, based on extensive Monte Carlo simulations of a simple off-lattice model, strongly support the view that knots play no role in enhancing a protein’s thermal stability or any other thermodynamic equilibrium property.
2 Model and Methods
2.1 The Cα Gō model
Proteins are represented by a simple Cα model (Figure 1). Accordingly, residues are reduced to hard spherical beads of uniform size, centered on the Cα atoms. Consecutive Cα atoms are connected by rigid sticks representing pseudobonds on the amide planes. We adopt a radius of 1.7 Å for the beads, which is the van der Waals radius of Cα atoms Tsai1999. For the length of each stick, we adopt the distance between the Cα atoms of the respective bonded residues in the protein’s native conformation, these being approximately 2.9 Å for cis bonds and 3.8 Å for trans bonds. Two non-bonded residues are said to be in contact in the native conformation if the smallest distance between any two heavy atoms, one belonging to each residue, is 4.5 Å, this cut-off being chosen because it is slightly larger than twice the average van der Waals radius of heavy atoms in proteins.
To model protein energetics, we consider the native-centric Gō potential Go. Accordingly, the total energy of a conformation defined by bead coordinates is given by
where is the chain length measured in the number of beads, is the position vector of bead in the native structure, is 1 if the contact is present in the native conformation and is 0 otherwise, is a uniform intramolecular energy parameter (taken as in this study, in which energies and temperatures are shown in reduced units), is the half-width of the inverse quadratic potential well, and the chirality of contact in the conformation under consideration is
The chirality of the contact in the native conformation is
In equations (2) and (3), is Heaviside’s unit step function, which takes the value 1 if its argument is greater than zero and the value 0 otherwise. The chirality factor in (1) favors the native conformation vis à vis its mirror conformation. A native contact is considered formed if the distance between the centers of the respective beads differs from the distance between their Cα atoms in the native conformation by less than the half-width of the potential wells, .
2.2 Replica-exchange Monte Carlo simulations
The conformational space of the C-alpha model is explored with Metropolis Monte Carlo (MC) Metropolis(1953) by means of a move set that comprises crankshaft and pivot moves. All simulations start from an unfolded (and unknotted) conformation obtained from a high-temperature simulation, and folding progress is monitored using the fraction of native contacts, . To sample equilibrium distributions, we use Monte Carlo replica-exchange (MC-RE) Sugita(1999).
To explore the conformational space, two sampling schemes can be deployed. A canonical one in which the move set preserves the linear topology of the chain, and another one that breaks the linear topology of the chain by allowing it to cross itself Especial2022, while preserving the sequential bonding of the protein chain and excluded volume among its units. We have previously shown, using simple Cα protein models (of unknotted proteins Fn-III and m, shallowly knotted protein MJ0366, and deeply knotted protein Rds3p), that there is no difference in equilibrium properties calculated with a move set that preserves linear topology and one that does not Especial2022. However, the move set that breaks the linear topology of the chain is extraordinarily advantageous in the case of deeply knotted proteins like the ones studied here, since it provides correct equilibrium results at a much lower computational cost Especial2022. Thus, unless otherwise stated, we will deploy the sampling method that does not preserve the linear topology of the chain.
The weighted histogram analysis method (WHAM) Chodera(2007) is used to analyze data from the MC-RE simulations and produce maximum likelihood estimates of the density of states, from which expected values for thermodynamic properties are calculated as functions of temperature. An example is the heat capacity, , defined in reduced units as . The melting temperature, , is determined as the temperature at which the peaks. We measure the cooperativity degree of the transition by the ratio of the full width at half maximum (FWHM) of the peak to the melting temperature, and the half-width of the potential well, , is adjusted to obtain a simulated FWHM ratio between 4 and 5% Larriva2010, Fernandez2021. This criterion has been successfully used in previous simulations employing a similar potential and sampling faisca2016, faisca2019. WHAM is additionally used to project the density of states along to obtain free energy profiles at a specific temperature.
2.3 Molecular Dynamics simulations
The MD simulations were performed using GROMACS 2024.3 abraham2015, with the AMBER 14SB force fields maier2015, and the TIP3P water model jorgensen1983, neria1996. The system consisted of a protein, 8808 water molecules, and 4 Cl- ions to equilibrate the overall system charges. We performed a two-step minimization procedure using the steepest descent method and an initialization protocol with two additional steps, in which temperature (NVT) and pressure (NPT) baths are turned on and equilibrated for 100 ps segments, with an integrator step of 1 fs. The nonbonded interactions were computed using a single cutoff of 1.0 nm, with the neighbor list updated every 20 steps. Van der Waals interactions were truncated beyond this cutoff, while long-range electrostatics were calculated using the Particle-Mesh Ewald (PME) method darden1993, with a Verlet cutoff of 1.0 nm and a Fourier grid spacing of 0.125 nm. Bond lengths involving hydrogen atoms in solute and water molecules were constrained using the LINCS hess2008b and SETTLE miyamoto1992 algorithms, respectively. In the NVT ensemble, the v-rescale thermostat bussi2007 was employed to maintain a temperature of 310 K (with a coupling constant of 0.1 ps). In the NPT ensemble, the v-rescale thermostat bussi2007 (coupling constant of 1.0 ps) was used in combination with the Parrinello-Rahman barostat parrinello1981 (coupling constant of 2.0 ps and isothermal compressibility of 4.510-5 bar-1). For the MD relaxation step, we run for 100 ns with the integration step adjusted to 2 fs.
2.4 Topological state
The topological state (knotted or unknotted) of a sampled conformation is determined using the Koniaris-Muthukumar-Taylor (KMT) algorithm taylor2000.
3 Results and discussion
3.1 Model systems
Proteins YbeA (from Escherichia coli, PDB ID: 1ns5) and YibK (from Haemophilus influenzae, PDB ID: 1mxi and PDB ID:1j85) are homodimeric enzymes that belong to the \-knot superfamily of MTases. In this work, we consider the monomer identified as chain A in the corresponding PDB files. The native structure of both proteins features a trefoil knot. In a knotted protein, the region of the polypeptide chain encompassing the knot is termed the knotted core. Protein knots are classified as deep or shallow, depending on the number of residues that must be removed from one of the termini to unknot the chain. If the latter is larger (or smaller) than 10, the knot is deep (or shallow). YbeA has a chain length of 155 amino acids, and its knotted core extends from residue 69 to residue 121. Consequently, it is necessary to remove 68 residues from the N-terminus (or 34 residues from the C-terminus) to untie the knot that is classified as deep. The native structure of YibK also embeds a deep trefoil knot. Indeed, YibK has a chain length of 160 residues and a knotted core that extends from residue 77 through residue 120, making it necessary to remove 76 residues from the N-terminus (or 40 residues from the C-terminus) to untie the knot. Despite their structural similarity, these bacterial proteins share only 19% of sequence identity.
To explore the role of the knotted topology in determining equilibrium properties, it is necessary to consider control systems Yeates2010. A suitable control system is a similar, yet unknotted protein that must be used for comparison. Since knotted/unknotted pairs do not exist in nature, they must be designed or engineered. In the present study, we consider two control systems obtained experimentally and one control system created through computational modeling. The in vitro control systems were engineered by Hsu and co-workers through a process known as circular permutation Danny2019a, Danny2019b. A circular permutant (CP) is an engineered protein that results from linking the C- and N-terminus of the polypeptide chain after disrupting the protein’s backbone at some selected site Faisca2012. In the case of protein YbeA, an unknotted control system was obtained by creating a CP after breaking the chain at the knotted core in posiiton 74, while for YibK the chain was broken at position 82, also in the knotted core, to create the CP (Figure 2A). In the present study, the resulting unknotted structures are designated by CP-YbeA (PDB ID: 5zyo) and CP-YibK (PDB ID: 6ahw). The in silico control structure was obtained for YibK (PDB ID: 1mxi) by manually passing the loop containing residues 114–125 across the knotted core (amino acids 81-85) (Figure 2B). The manipulated structure was then subjected to an MD simulation, in which segments 1-80, 86-113, and 126-156 were positionally restrained 1000 kJ/mol nm2), while the two loops (81–85 and 114–125) were allowed to relax. The final relaxed and refined structure is named IS-YibK.
For all three structures, we computed the gyration radius (), which is a measure of structural compactness, the root-mean-square-deviation (RMSD) to the wild-type native structure, which is a measure of structural similarity, and the number of native contacts, which is a measure of the system’s energy in the context of the adopted model. These properties are reported in Table 1. The native structure of CP-YbeA exhibits significant structural rearrangements, which are visible in the cartoon representation (Figure 2A), and result in a substantially high RMSD (3Å). There is also a significant loss of compactness, resulting in a high gyration radius and a substantial loss (25%) of native contacts. On the other hand, both the unknotted CP-YibK and IS-YibK share a high degree of structural similarity with YibK (RMSD2Å). However, while IS-YibK has fewer native contacts than YibK (11%), the circular permutant exhibits a nearly identical degree of compactness, and essentially the same number of native contacts (the difference between the two model systems being only 2%), which makes it the best control system considered here.
For each knotted protein and corresponding control system, we measured several equilibrium properties as a function of temperature.
| Protein | Topological state | RMSD (Å) | (Å) | No. native contacts |
| YbeA | K | - | 15.5 | 542 |
| CP-YbeA | U | 16.6 | 21.8 | 410 |
| YbiK | K | - | 15.9 | 552 |
| CP-YibK | U | 0.35 | 15.5 | 561 |
| IS-YibK | U | 0.48 | 16.6 | 500 |
3.2 Protein YbeA and its control system
Results reported in Figure 3 (A-D) show that YbeA folds cooperatively in a two-state manner, while the control CP-YbeA folds non-cooperatively. This lack of cooperativity is particularly evident from the rather gentle slope of the internal energy curve (Figure 3 A), and from the broadness of the heat capacity curve (Figure 3 D). In addition, the melting temperature of CP-YbeA is significantly (15%) lower than that of YbeA. We have also measured the dependence of the free energy on the fraction of native contacts () at the melting temperature for the knotted-unknotted pair. We find that the behavior of the folding transition changes from two-state (for YbeA) to downhill (or barrierless) (for CP-YbeA) (Figure 4 A). These findings, in particular the decreased thermal stability of CP-YbeA, cannot be attributed solely to its unknotted topological state. Indeed, untying the knot in YbeA introduced drastic structural changes and a significant loss of native interactions, which collectively account for the observed behavior.
3.3 Protein YibK and its control systems
To isolate the role of the topological state on thermal equilibrium properties, particularly thermal stability, we now examine YibK and its control systems. In this case, the two-state nature of the folding transition is conserved across the three model systems (Figure 3E-H). A minor, yet non-negligible difference is observed between YibK and its in silico control system, IS-YibK. In particular, for the latter, the melting temperature is 7% lower than that of YibK (Figure 3H). Since the number of native contacts is not conserved in the unknotted protein, we cannot ascribe its lower thermal stability to its unknotted topological state. In fact, when one compares YibK and CP-YibK, which are nearly structurally identical and share almost the same number of native contacts, it is found that their thermodynamic behavior is practically coincident, with the thermal stability differing by only 1%. In line with these findings, the similarity between the free energy profiles (Figure 4B) is also significantly higher for YibK and its unknotted circular permutant.
3.4 Comparison with experimental data
The results reported in the previous section show that as the similarity between the knotted protein and its unknotted control increases, the differences in the thermodynamic equilibrium behavior tend to disappear. In particular, they indicate that if one could construct an unknotted control system conserving the same native geometry (i.e., the atomic positions and, consequently, distances between atoms), one should observe the same thermodynamic behavior. These findings support the view that the protein’s topological state does not enhance thermal stability or affect thermodynamic equilibrium properties. From a physics standpoint, this result is expected since the protein’s energy is not a function of the topological state, and, consequently, the equilibrium density of states should be neither, provided the exploration of the conformational space is ergodic. An identical density of states is expected for both knotted and unknotted control systems, which translates into identical thermodynamic behavior.
However, the experimental measurements carried out for YibK and its unknotted control system contradict this argument, indicating that the knotted topology enhances the protein’s thermal stability. Indeed, a decrease in the melting temperature of 22% (measured with differential scanning calorimetry) or 27% (measured with far-UV Circular Dichroism (CD) signals) was reported Danny2019a for the unknotted protein.
Part of the decrease in thermal stability observed experimentally may be due to the fact that the dimer interface for each chain in the dimeric version of CP-YibK is reduced by 13%, which leads to a loss of intermolecular hydrogen bonds and salt bridges near the dimer interface Danny2019a. Since experimental data refers to the dimeric form of the protein, and simulations focused on the monomer, a direct comparison between experimental and computational data is not possible. That said, it is clear that the experimental and computational results offer opposite views on the role of topology in setting YibK’s thermal stability.
A few years ago, Jennings and co-workers reported two distinct timescales for the unfolding and the untying of protein MTTTm (PDB ID: 106d) – a protein of the same family as YibK that features a similar chain length. Their findings showed that the transition to untied states required at least 6 months to occur Jennings2016, while the transition to unfolded states occurs in 2 weeks. This slow unknotting process suggests that, for deeply knotted proteins like YibK, the full unfolding–untying transition may not be observed within the timescale of a DSC experiment, indicating a deviation from equilibrium thermodynamics.
These experimental observations align with insights from computational simulations. As previously noted, it is particularly challenging to achieve equilibrium distributions for complex polymeric structures — such as deeply knotted proteins — using MC simulations based on a move set that preserves the linear topology of the chain, and starting from unfolded and unknotted conformations. For example, in the case of protein Rds3p, which is 51 amino acids shorter than YibK, simulations preserving linear topology required at least an order of magnitude more MC steps to reach equilibrium compared to simulations using move sets that allow the chain to cross itself Especial2022. Notably, for unknotted proteins (e.g., Fn-III and beta-2-microglobulin), the computational cost is independent of the move set Especial2022.
Given that real protein dynamics also preserve the linear topology of the polypeptide chain, it is plausible that the scan rate (i.e., heating rate) used in the DSC experiment was too fast for the system to equilibrate at each temperature, and even more so to allow the population of unknotted conformations. This supports the view that the experimental measurements likely reflect a non-equilibrium distribution lacking unfolded and unknotted conformations, and, consequently, the measured melting temperature represents an upper bound for the real melting temperature.
4 Conclusions
Over the past 20 years, a remarkably large body of theoretical and experimental work has been dedicated to the study of knotted proteins. A major line of research has been devoted to understanding whether knots may confer any functional advantage to their carriers Jackson2020. Answering this question will provide insight into the evolutionary significance of knotted proteins and may have a potential impact on designing drugs for technological or medical applications.
A major challenge in studying the role of knots in proteins is the construction of a suitable control system, a version of the protein that is structurally similar to the knotted variant but lacks the knot. In the context of lattice models, it is quite straightforward to untie a model protein while introducing minimal changes to the system; previous simulation efforts in the lattice context have shown that a knotted topology does not enhance the carrier’s thermal stability Soler2013. Unfortunately, in real-world proteins, this is not as simple. Here, we employed a simple off-lattice model to show that as the structural similarity between the protein and its control system increases, the difference in thermodynamic equilibrium properties vanishes. In particular, we have provided compelling evidence that the knot in protein YibK does not enhance its thermal stability. This result is in contradiction with recent experimental data on this protein, which reported a moderate increase in thermal stability for the knotted version of Yibk Danny2019a.
Considering that the timescale required for untying a protein knot can be orders of magnitude longer than that for unfolding Jennings2016, and that Monte Carlo simulations face significant challenges in reaching equilibrium distributions for deeply knotted proteins like YibK when starting from unknotted conformations, we propose that the experimental data may reflect an unfolding–folding transition occurring under non-equilibrium conditions.
5 Author contributions
PFNF designed the research, prepared the figures, and wrote the manuscript. JNCE developed the code and performed all the calculations. All authors analyzed the data. AN, BPT, and MM prepared the in silico unknotted version of YibK.
6 Declaration of interests
The authors declare no competing interests.
Work supported by UID/04046/2025 Instituto de Biosistemas e Ciências Integrativas Centre Grant from Instituto de Biosistemas e Ciências Integrativas Centre grant from FCT, Portugal. J.N.C.E. acknowledges financial support from FCT, Portugal, through PhD grant SFRH/BD/144345/2019. B.P.T. acknowledges financial support from Fundação Calouste Gulbenkian through the Gulbenkian Novos Talentos Fellowship. This study was also supported by the European Union (TWIN2PIPSA - Twinning for Excellence in Biophysics of Protein Interactions and Self-Assembly, GA 101079147). A part of this work was performed on the computational resources of INCD (http://www.incd.pt) funded by FCT and UE under project LISBOA-01-0145-FEDER-022153. Access was granted by FCT through project 2022.26279.CPCA.A0.